A psychologist has developed an aptitude test which consists of a series of mathematical and vocabulary problems. They want to test the hypothesis that the mean test score is 64.6.

A random sample of 40 people have taken the test and their results recorded:

Download the data

69 45 62 71 67 72 68 62 69 64
81 57 95 85 56 60 76 82 61 55
75 48 93 74 63 98 74 52 72 81
44 79 57 54 44 69 55 91 52 88

You may find this Student's t distribution table useful throughout this question.

a)Calculate the test statistic (t) for the hypothesis test. Give your answer to 4 decimal places.

t =

b)A level of significance of α = 0.05 is to be used for the test. The P-value for this test statistic is less, or greater than the level of significance.

c)The result of this test is that the null hypothesis is rejected, or not rejected.
Answer #1
Values ( X ) Σ ( Xi- X̅ )2
69 1
45 529
62 36
71 9
67 1
72 16
68 0
62 36
69 1
64 16
81 169
57 121
95.0 729
85 289
56 144
60 64
76 64
82 196
61 49
55 169
75 49
48 400
93 625
74 36
63 25
98 900
74 36
52 256
72 16
81 169
44 576
79 121
57 121
54 196
44 576
69 1
55 169
91 529
52 256
88 400
Total 2720 8096

Mean X̅ = Σ Xi / n
X̅ = 2720 / 40 = 68
Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 8096 / 40 -1 ) = 14.408

To Test :-
H0 :- µ = 64.6
H1 :- µ ≠ 64.6

Test Statistic :-
t = ( X̅ - µ ) / ( S / √(n))
t = ( 68 - 64.6 ) / ( 14.408 / √(40) )
t = 1.4925


Test Criteria :-
Reject null hypothesis if | t | > t(α/2, n-1)
t(α/2, n-1) = t(0.05 /2, 40-1) = 2.023
| t | > t(α/2, n-1) = 1.4925 < 2.023
Result :- Fail to reject null hypothesis

Decision based on P value
P - value = P ( t > 1.4925 ) = 0.1436
Reject null hypothesis if P value < α = 0.05 level of significance
P - value = 0.1436 > 0.05 ,hence we fail to reject null hypothesis
Conclusion :- Fail to reject null hypothesis